OpenStudy (anonymous):

find x and show working 6+2^x=192

OpenStudy (anonymous):

6+2^x=192 2^x=186 xln(2)=ln(186) x=[ln(186)]/[ln(2)]

OpenStudy (anonymous):

2^x=186